OpenStudy (anonymous):

4, 15, 10, 8, 33, 25, 36, 66, 63, 50, 64, 49, 40. Find variance?

OpenStudy (anonymous):

I can get the answer from online calculator but I need to know the steps...

OpenStudy (anonymous):

Can you find mean here??

OpenStudy (anonymous):

You just want square of mean deviations here.. Firstly find Mean and then mean deviation..

OpenStudy (anonymous):

Okay.. Simply add all the data here.. @pagalbilal

OpenStudy (anonymous):

mean is 35.6154

OpenStudy (anonymous):

Yes Mean is right..

OpenStudy (anonymous):

You have to do long calculation here.. Now: |4 - 35.6| = ?? |15 - 35.6| = ?? |10 - 35.6| = ?? Like wise subtract all the data from mean and you will get here Mean Deviations..

OpenStudy (anonymous):

ok what about "The number of complaint letters received by the USPS in a given day are' is that quantitative or qualitative ?

OpenStudy (anonymous):

ohhh I see...I m working on it nw

OpenStudy (anonymous):

Then take the square of every difference also and add them like: for first one : |4 - 35.6| = 31.6 (now square it : 998.56) |15 - 35.6| = 20.6 (Now square it: 424.36) Like wise finding and squaring all the differences then you will add them.. By this you will get Standard Deviation but you are to find Variance: As Variance is square of Standard Deviation... After finding SD then square it you will get Variance..

OpenStudy (anonymous):

ok thank you very much for your help...